during an impact time lasting approximately 5.0 × 10^4s a golf club exerts an average force of about 5000N on the golf ball.what linear impulse is delivered to the ball as a result?
1
Expert's answer
2020-10-19T13:22:05-0400
"\Delta I=F\Delta t\\\Delta I=(5\cdot 10^4)(5000)=2.5\cdot 10^8\ Ns"